# The InSAR absolute phase amid singularities

### v1

@@ Abstract
The radar interferometric absolute phase is essential for estimating topography and displacements. However, its conventional definition based on the range difference is idealized in that it cannot be applied to complex, dynamic targets. Here, a universal observational definition is proposed, which is easiest to describe for differential interferometry: The absolute phase is determined by temporally unwrapping the phase while continuously varying the intermediate acquisition time between primary and secondary acquisitions. This absolute phase is typically not directly observable because a continuous series of observations is required. The absolute phase of a point target is proportional to the range difference, matching the conventional definition. For general targets undergoing a cyclic change, the absolute phase may be nonzero and then cannot be interpreted as a range difference. When a phase singularity (vanishing coherence) occurs at an intermediate time, the absolute phase becomes undefined, a situation termed an absolute phase singularity. Absolute phase singularities complicate absolute phase reconstruction through multifrequency techniques and through unwrapping multidimensional interferograms. They leave no trace in an interferogram, but unwrapping paths need to avoid those across which the absolute phase jumps by nonzero integer multiples of $2\pi$. Mathematical analyses identify conditions for unwrapping-based reconstruction up to a constant, accounting for absolute phase singularities, undersampling and noise. The general definition of the absolute phase and the mathematical analyses enable a comprehensive appraisal of InSAR processing chains and support the interpretation of observations whenever low coherence engenders phase and absolute phase singularities.
@@ 1 Introduction
In radar interferometry, the absolute phase is an essential but nevertheless ill-defined concept [1, 2]. According to Madsen et al. [3], the absolute phase is proportional to the difference in range between the primary and secondary acquisition from which the interferogram was formed. A similar view is adopted in optical interferometry [4]. While the absolute interferometric phase is not observed directly, many authors assert that it needs to be known for estimating topography or displacements [5, 6, 2]. But a general definition of the absolute phase is lacking. What is the absolute phase over a time period in spring in which snow melts and leaves flush? To answer this and similar questions, Sec. 23 establish an observational definition of the absolute phase in terms of measurable quantities. This definition is applicable to any InSAR (Interferometric Synthetic Aperture) mode and does not require any knowledge about how the electromagnetic wave interacts with the observed target.
The observational definition of the absolute phase $\phi^{\mathrm{a}}(\mathbf{w}_{2})$ postulates a continuous series of interferometric phase measurements. Throughout, the primary configuration $\mathbf{w}_{1}$, including the antenna position and time, is held fixed. The secondary configuration $\mathbf{w}$ is continuously varied along a pre-specified absolute path that starts at $\mathbf{w}_{1}$ and ends at the actual secondary configuration $\mathbf{w}_{2}$. Unwrapping the interferometric phase as the secondary configuration is varied [7] produces the absolute phase. For differential InSAR with a temporal baseline, the absolute phase is determined by unwrapping the interferometric phase over time. The observational and the conventional definition agree for an ideal point target.
@@
Fig. 1 shows the absolute phase $\phi^{\mathrm{a}}(s)$ modeled for three moving surface segments as a function of dimensionless time $s$ for a fixed primary at $s=0$. For $s<1/3$, segments 2 and 3 move up simultaneously, pulling the phase to $-2\pi$. As the segments return to their initial position at secondary time $s=1$, the absolute phase $\phi^{\mathrm{a}}(1)=-2\pi$. This nonzero value cannot be interpreted as a range difference, as the positions are identical at $s=0$ and $s=1$. Had the segments remained stationary, the observed wrapped phase would have been zero for all $s$, and $\phi^{\mathrm{a}}(1)=0$. Such path dependence and related characteristics will be studied in examples from differential InSAR and across-track InSAR modes in Sec. 4 and 5.
Figure 1: The differential InSAR absolute phase $\phi^{\mathrm{a}}(s)$ as a function of dimensionless time $s$ for three surface segments. Their normalized positions $b_{1}$, $b_{2}$, and $b_{3}$ are shown in the three bottom panels and also indicated in the insets. Model details can be found in Sec. 4.3.2
Path dependence controls how the absolute phase behaves across absolute phase singularities. In Sec. 6, it will be shown that the absolute phase can be patched continuously across absolute phase singularities if and only if there is no path dependence, provided certain continuity conditions hold. Using ideas from algebraic topology [8], we will see how path dependence arises and, ultimately, how it complicates the absolute phase reconstruction from actual observations.
@@ 2.1.1 Definition
The wrapped interferometric phase $\varphi$ is derived from two individual complex signals $u_{1}$ and $u_{2}$ by cross-multiplication [1], viz.
@@
[…]
@@
The wrapped phase $\varphi=\arg(\gamma)$ is only defined for $\rho>0$, while a phase singularity occurs for $\rho=0$. The engineering sign convention adopted throughout requires a positive sign in the complex exponential in (5). To convert to the physics sign convention with a negative sign, $\mathsf{j}$ needs to be replaced by $-\mathsf{i}$. An idealized measurement (e.g., no oscillator noise) produces error-free phase observations, as will be assumed in the theoretical analyses. This implies $\varphi=0$ for identical primary and secondary signals. There is thus no gauge freedom in the interferometric phase [9].
@@ 2.1.2 Mathematical representation
The wrapped phase $\varphi$ is naturally thought of as a point on the unit circle $S^{1}$ in the complex plane [8]. The argument mapping $\arg:\mathbb{\mathbb{C}^{+}}\rightarrow S^{1}$ is a radial projection from the non-zero complex number $\gamma\in\mathbb{C}^{+}$ onto $S^{1}$. The circular nature arises from the periodicity of the complex exponential, as $\exp\mathsf{j}\varphi=\exp\mathsf{j}\left(\varphi+2\pi N\right)$ for any $N\in\mathbb{Z}$. We can capture this periodicity by identifying the wrapped phase with an equivalence class [10]
$[\varphi]=\{\varphi+2\pi N\,\big{|}\,N\in\mathbb{Z}\}\,,$ (6)
an element of the quotient $\mathbb{R}/(2\pi\mathbb{Z})$.
Representing the wrapped phase by a number in an interval is problematic because the circle wraps back on itself. The canonical representation in the half-open interval $(-\pi,\pi]$ introduces an arbitrary discontinuity along the negative real axis [11]. We could also choose a different interval such as $(-2\pi,0]$, corresponding to a different discontinuity or, in complex analysis language, a different branch cut of the complex logarithm.
@@ 2.1.3 Dependence on variables
The coherence, and in turn $\varphi$, commonly depend on exogenic variables. For defining the absolute phase, the exogenic variables describe InSAR configuration parameters such as the observation time and antenna position of the secondary acquisition, while the primary configuration is fixed. The configuration parameters are encoded by $\mathbf{w}$ in a space $W$, assumed to be a subset of $R^{n}$. For modeling, it can be convenient to fix the configuration parameters and introduce model parameters $\mathbf{b}$ in a space $B$; they may for instance describe a scatterer’s position. For describing observed InSAR coherences, the coherence will vary with a space-like variable $\mathbf{x}\in X$, with e.g. $X=\mathbb{R}^{2}$ for a planar interferogram. For now, we will focus on a configuration parameter space $W$.
@@
The wrapped phase depends on $\mathbf{w}$ as $\varphi(\mathbf{w})=\arg(\gamma(\mathbf{w}))$. Its domain excludes the phase singularities, i.e.
@@
where the pre-image of 0 is denoted by $\gamma^{-1}\left[\{0\}\right]$. In terms of functions, we have a mapping $\varphi:W^{+}\rightarrow S^{1}$ given by the composition $\varphi=\arg\circ\gamma$, depicted graphically in Fig. 2a). (To be precise, the function $\gamma$ needs to be restricted to $W^{+}$.) The wrapped phase $\varphi$ inherits the $C^{\infty}$ smoothness from $\gamma$.
@@ 2.1.4 Phase singularities
InSAR phase singularities are zeros of a coherence function, widely studied in optics [12, 11]. Phase singularities in zeros of complex scalar wavefields in e.g. optics and acoustics and of quantum mechanical wavefunctions, are also called optical vortices or wavefront dislocations; they feature in such phenomena as the Goos-Hänchen shift, seiches on a rotating Earth, or the ocean tides [13, 14, 15]. For an overview of their history and properties, see [16, 11, 15]. Their topological characteristics, for instance how they form and annihilate in bifurcations when the complex field is disturbed, are discussed in [17, 18, 19].
Phase singularities generically occur as points in the plane ($\mathbb{R}^{2}$). That is, almost all phase singularities are points [16], as that in Fig. 2. Non-point singularities may arise from symmetry, such as the circles of vanishing amplitude in a circular aperture’s Fraunhofer diffraction pattern. In $\mathbb{R}^{3}$, generic phase singularities form curves, including lines and loops [20, 18]. In a general space $W$, their dimension is two less than that of $W$. Their codimension is two [11], as both the real and imaginary part need to vanish.
A key characteristic of a phase singularity is that the phase can spiral around it [15], as in Fig. 2b. Along a loop around a phase singularity, the phase changes by an integer multiple of $2\pi$, this change being an unwrapped phase difference.
@@ 2.2.1 Definition
An unwrapped phase difference $\Delta\phi\in\mathbb{R}$ is an unambiguous measure of the difference between wrapped phases at two points connected by a path in a space $W$. To obtain that difference between a wrapped $\varphi(\mathbf{w})\in S^{1}$ at a point $\mathbf{w}\in W^{+}$ and another wrapped phase $\varphi(\mathbf{w}^{\prime})$ at $\mathbf{w}^{\prime}\in W^{+}$, unwrapping [7] integrates phase increments along a path connecting $\mathbf{w}$ with $\mathbf{w}^{\prime}$. The result is not subject to the $2\pi N$ ambiguity of the wrapped phase. Intuitively, this is because two sufficiently close points $\mathbf{w}$ and $\mathbf{w}+\mathrm{d}\mathbf{w}$ will have almost the same phase due to the continuity of $\varphi(\mathbf{w})$, effectively removing the ambiguity [8]. The integration accumulates these unambiguous increments.
For this unwrapping process, a connecting path $\Gamma$ from $\mathbf{w}^{\prime}$ to $\mathbf{w}$ has to be specified. The rectifiable path $\Gamma$ is taken by a continuous, piecewise smooth function from the unit interval $I\equiv[0,1]$ to $W^{+}$ such that $\Gamma(0)=\mathbf{w}^{\prime}$ and $\Gamma(1)=\mathbf{w}$.
Mathematically, the unwrapped phase difference $\Delta\phi(\Gamma)$ is given by [7]
[…]
where the differential one-form $d\varphi_{W^{+}}(\mathrm{d}\mathbf{w})$ determines the infinitesimal phase change for an infinitesimal step $\mathrm{d}\mathbf{w}$. The one-form is often expressed as $\nabla\phi\cdot\mathrm{d}\mathbf{w}$, with $\nabla\phi$ referred to as a gradient. However, in general (8) depends on the path $\Gamma$, while it would not if $\nabla\phi$ were a true gradient. In the language of differential forms, $d\varphi_{W^{+}}$ is, in general, not exact [21].
@@ 2.2.2 Evaluation
To evaluate the integral, we map the path $\Gamma$ into the punctured complex plane $\mathbb{C}^{+}=C\setminus{0}$. This is done through the coherence mapping $\gamma$ (restricted to $W^{+}$), or alternatively through $v$. Representing the coherence as $\gamma=\xi+\mathsf{j}\eta$, the change in phase relates to changes in $\xi$ and $\eta$ as

### v3

@@ Abstract
The radar interferometric absolute phase is considered essential for estimating topography and displacements. However, the conventional definition—that absolute phase is proportional to the range difference—cannot be applied to general targets. Here, a universal observational definition is proposed, which is easiest to describe for differential interferometry: The absolute phase is determined by temporally unwrapping the phase as the intermediate secondary acquisition time sweeps from the primary to the secondary acquisition time. This absolute phase is typically not directly observable because a continuous series of observations is required. The absolute phase of a point target is proportional to the range difference, matching the conventional definition. For general targets undergoing cyclic change (returning to their initial state), the absolute phase may be nonzero and cannot be interpreted as a range difference. When a phase singularity (vanishing coherence) occurs at an intermediate time, the absolute phase becomes undefined, a situation termed an absolute phase singularity. Absolute phase singularities complicate absolute phase reconstruction through multifrequency techniques and through unwrapping multidimensional interferograms. They leave no trace in an interferogram, but unwrapping paths need to avoid those across which the absolute phase jumps by nonzero integer multiples of $2\pi$. Mathematical analyses identify conditions for unwrapping-based reconstruction up to a constant, accounting for absolute phase singularities, undersampling and noise. The general definition of the absolute phase and the mathematical analyses enable a comprehensive appraisal of InSAR processing chains and support the interpretation of observations whenever low coherence engenders phase and absolute phase singularities.
@@ 1 Introduction
In radar interferometry, the absolute phase is an essential but nevertheless ill-defined concept [bamler98, rosen00]. According to Madsen et al. [madsen93], the absolute phase is proportional to the range difference between the primary and secondary acquisition from which the interferogram was formed. A similar view is adopted in optical interferometry [creath85]. While the absolute interferometric phase is not observed directly, many authors assert that it needs to be known for estimating topography or displacements [lombardini98, zebker98, rosen00]. But a general definition of the absolute phase is lacking. What is the absolute phase over a time period in spring in which snow melts and leaves flush, while coherence is partially retained? To answer this and similar questions, Sec. 23 establish an observational definition of the absolute phase in terms of measurable quantities. This definition is applicable to any InSAR (Interferometric Synthetic Aperture) mode and does not require any knowledge about how the electromagnetic wave interacts with the observed target.
[…]
@@
An example of path dependence is shown in Fig. 1, depicting the absolute phase $\phi^{\mathrm{a}}(s)$ modeled for three moving surface segments as a function of dimensionless time $s$ for a fixed primary at $s=0$. The surface segments, not resolved in the observations, are inspired by permafrost landforms comprised of multiple parts that undergo wavelength-scale differential movements [zwieback24]. Segment 1 (red) stays stationary throughout, at position $b_{1}(s)=0$. Conversely, segments 2 (teal) and 3 (blue) move up simultaneously for $s<1/3$. Upward movement corresponds to a decreasing absolute phase, obtained by unwrapping over secondary time $s$. At secondary time $s=1$, all segments have returned to their initial position, but the absolute phase $\phi^{\mathrm{a}}(1)=-2\pi$. This nonzero value cannot be interpreted as a displacement or range difference, as the positions are identical at $s=0$ and $s=1$. Had the segments remained stationary, the absolute phase $\phi^{\mathrm{a}}(s)=0$ for all $s$.
Figure 1: The differential InSAR absolute phase $\phi^{\mathrm{a}}(s)$ as a function of dimensionless time $s$ for three surface segments. The InSAR observations do not resolve the segments and are continuous in $s$. The segments’ normalized positions $b_{1}$, $b_{2}$, and $b_{3}$ are shown in the three bottom panels and also indicated in the insets. Model details can be found in Sec. 4.3.2.
The loss of correspondence between absolute phase and range difference may seem undesirable, but it is not a quirk of the observational definition of the absolute phase. In differential interferometry the observationally defined absolute phase is the unique real-valued quantity that depends continuously on secondary time and is congruent with the wrapped phase (Sec. 4). In other words, proportionality to range difference is incompatible with a real-valued phase when there is path dependence. Such path dependence and related characteristics will be studied in examples from differential InSAR and across-track InSAR modes in Sec. 4 and 5.
The path dependence exhibited by complex targets controls how the absolute phase behaves across absolute phase singularities. In Sec. 6, it will be shown that the absolute phase can be patched continuously across absolute phase singularities if and only if there is no path dependence, provided certain continuity conditions hold. Using ideas from algebraic topology [fulton97], we will see how path dependence arises and, ultimately, how it complicates absolute phase reconstruction from actual observations.
@@ 2.1.1 Definition
The wrapped interferometric phase $\varphi$ is derived from two individual complex signals $u_{1}$ and $u_{2}$ by cross-multiplication [bamler98], viz.
@@
The angular brackets $\langle\cdot\rangle$ denote an appropriate averaging operation, including an ensemble average (for modeling), a spatial multilooking operation, or no averaging (for single-look interferometry). I adopt the convention that $u_{1}$ – the linear term – is the primary signal and $u_{2}$ – the conjugate-linear term involving complex conjugation – is the secondary signal.
@@
The wrapped phase $\varphi=\arg(\gamma)$ is only defined for $\rho>0$, while a phase singularity occurs for $\rho=0$. The engineering sign convention adopted throughout requires a positive sign in the complex exponential in (5). To convert to the physics sign convention with a negative sign, $\mathsf{j}$ needs to be replaced by $\mathsf{i}$. An idealized measurement (e.g., no oscillator noise) produces error-free phase observations, as will be assumed in the theoretical analyses. This implies $\varphi=0$ for identical primary and secondary signals. There is thus no gauge freedom in the interferometric phase [dennis01].
@@ 2.1.2 Mathematical representation
The wrapped phase is commonly represented by a number in an interval, but this introduces an arbitrary discontinuity in the complex plane. For the canonical representation in the half-open interval $(-\pi,\pi]$, the wrapped phase jumps by $\pm 2\pi$ across the negative real axis [dennis09]. Conversely, if we chose a different interval such as $(-2\pi,0]$, the discontinuity on the negative real axis would disappear, to be replaced by a different discontinuity or, in complex analysis language, a different branch cut of the complex logarithm.
The wrapped phase $\varphi$ is naturally thought of as a point on the unit circle $S^{1}$ in the complex plane [fulton97]. A $\varphi\in S^{1}$ is obtained by radially projecting a non-zero complex number $\gamma\in\mathbb{C}^{+}$ onto $S^{1}$, through the argument mapping $\arg:\mathbb{\mathbb{C}^{+}}\rightarrow S^{1}$. The circular nature arises from the periodicity of the complex exponential, as $\exp\mathsf{j}\tilde{\varphi}=\exp\mathsf{j}\left(\tilde{\varphi}+2\pi N\right)$ for any $N\in\mathbb{Z}$ and any $\tilde{\varphi}\in\mathbb{R}$.
An alternative representation capturing the circular nature uses equivalence classes. A $\tilde{\varphi}\in\mathbb{R}$ represents an equivalence class $\tilde{\varphi}$] of elements $\tilde{\varphi}^{\prime}\in\mathbb{R}$ with $\exp\mathsf{j}\tilde{\varphi}=\exp\mathsf{j}\tilde{\varphi}^{\prime}$, given by
$[\tilde{\varphi}]=\{\tilde{\varphi}+2\pi N\,\big|\,N\in\mathbb{Z}\}\,.$ (6)
In abstract terms, such an equivalence class is an element of the quotient $\mathbb{R}/(2\pi\mathbb{Z})$.
@@ 2.1.3 Dependence on variables
The coherence, and in turn $\varphi$, commonly depend on variables. For defining the absolute phase, these variables describe InSAR configuration parameters such as the observation time and antenna position of the secondary acquisition, while the primary configuration is fixed. The configuration parameters are encoded by $\mathbf{w}$ in a space $W$, assumed to be a subset of $\mathbb{R}^{n}$. For modeling, it can be convenient to fix the configuration parameters and introduce model parameters $\mathbf{b}$ in a space $B$; they may for instance describe a scatterer’s position. For describing observations such as a planar interferogram, we use $\mathbf{x}\in X$ to denote spatial variables; for a two-dimensional interferogram, the space $X$ is $\mathbb{R}^{2}$. For now, we will focus on a configuration parameter space $W$.
@@
The wrapped phase depends on $\mathbf{w}$ as $\varphi(\mathbf{w})=\arg(\gamma(\mathbf{w}))$. This expression does not make sense for phase singularities. We thus define the domain of $\varphi(\mathbf{w})$ as
@@
where the set difference $\setminus$ removes the pre-image of 0, denoted by $\gamma^{-1}\left[\{0\}\right]$. A point $\mathbf{w}\in W^{+}$ is in the pre-image of zero precisely if its image $\gamma(\mathbf{w})=0$, or in other words, precisely if it is a phase singularity. In terms of functions, we have a mapping $\varphi:W^{+}\rightarrow S^{1}$ given by the composition $\varphi=\arg\circ\gamma$. Explicitly, $\varphi(\mathbf{w})=(\arg\circ\gamma)(\mathbf{w})=\arg\left(\gamma\left(\mathbf{w}\right)\right)$, which is depicted graphically by the commutative diagram in Fig. 2a). (The function $\gamma$ is restricted to $W^{+}$.) The wrapped phase $\varphi$ inherits the $C^{\infty}$ smoothness from $\gamma$.
@@ 2.1.4 Phase singularities
Phase singularities as zeros of $\mathbb{C}$-valued functions are ubiquitous in physics. They commonly occur as zeros of coherence functions (e.g., InSAR, optics) and as zeros of complex scalar wavefields (e.g., optics, acoustics, quantum mechanical wavefunctions). The latter are also called optical vortices or wavefront dislocations and feature in such phenomena as the Goos-Hänchen shift, seiches on a rotating Earth, or the ocean tides [nye74, mortimer76, berry23]. For an overview of their history and properties, see [berry09, dennis09, berry23]. Their topological characteristics, for instance how they form and annihilate in bifurcations when the complex field is disturbed, are discussed in [freund95, berry06, adachi07].
Phase singularities generically occur as points in the plane ($\mathbb{R}^{2}$). That is, almost all phase singularities are points [berry09], as that in Fig. 2. Non-point singularities may arise from symmetry, such as the circles of vanishing amplitude in a circular aperture’s Fraunhofer diffraction pattern. In $\mathbb{R}^{3}$, generic phase singularities form curves, including lines and loops [huntley01, berry06]. In a general space $W$, their dimension is two less than that of $W$. Their codimension is two [dennis09], as both the real and imaginary part need to vanish.
A key characteristic of a phase singularity is that the phase can spiral around it [berry23], as in Fig. 2b. Along a loop around a phase singularity, the phase changes by an integer multiple of $2\pi$, this change being an unwrapped phase difference.
@@ 2.2.1 Definition
An unwrapped phase difference $\Delta\phi\in\mathbb{R}$ is an unambiguous measure of the difference between wrapped phases at two points connected by a path in a space $W$. To obtain that difference between a wrapped $\varphi(\mathbf{w})\in S^{1}$ at a point $\mathbf{w}\in W^{+}$ and another wrapped phase $\varphi(\mathbf{w}^{\prime})\in S^{1}$ at $\mathbf{w}^{\prime}\in W^{+}$, unwrapping [ghiglia98] integrates phase increments along a path connecting $\mathbf{w}$ with $\mathbf{w}^{\prime}$. The result is not subject to the $2\pi N$ ambiguity of the wrapped phase. Intuitively, this is because two sufficiently close points $\mathbf{w}$ and $\mathbf{w}+\mathrm{d}\mathbf{w}$ will have almost the same phase due to the continuity of $\varphi(\mathbf{w})$, effectively removing the ambiguity [fulton97]. The integration accumulates these unambiguous increments.
For this unwrapping process, a connecting path $\Gamma$ from $\mathbf{w}^{\prime}$ to $\mathbf{w}$ has to be specified. The rectifiable path $\Gamma$ is given by a continuous, piecewise smooth function from the unit interval $I\equiv[0,1]$ to $W^{+}$ such that $\Gamma(0)=\mathbf{w}^{\prime}$ and $\Gamma(1)=\mathbf{w}$.
Mathematically, the unwrapped phase difference $\Delta\phi(\Gamma)$ is given by [ghiglia98]
$\Delta\phi(\Gamma)=\int_{\Gamma}d\varphi_{W^{+}}\,$, (8)
where $d\varphi_{W^{+}}$ is the differential one-form that measures phase changes. As a one-form, it takes an infinitesimal step $\mathrm{d}\mathbf{w}$ as input and returns the infinitesimal phase change corresponding to $\mathrm{d}\mathbf{w}$. For our purposes, it is sufficient to think of a differential $k$-form as an integrand in a $k$-dimensional integral [weintraub14]. The one-form $d\varphi_{W^{+}}$ is integrated along a (one-dimensional) path, such as $\Gamma$ or $\tilde{\Gamma}$ shown in Fig. 2b.
The one-form $d\varphi_{W^{+}}$ is often expressed as $\nabla\phi\cdot\mathrm{d}\mathbf{w}$, with $\nabla\phi$ referred to as a gradient. However, in general (8) depends on the path $\Gamma$, while it would not if $\nabla\phi$ were a true gradient. In the language of differential forms, $d\varphi_{W^{+}}$ is, in general, not exact [weintraub14].
@@ 2.2.2 Evaluation
To evaluate the integral, we map the path $\Gamma$ into the punctured complex plane $\mathbb{C}^{+}=C\setminus{0}$. This is done through the coherence mapping $\gamma$, or alternatively through $v$. Representing the coherence as $\gamma=\xi+\mathsf{j}\eta$, the change in phase relates to changes in $\xi$ and $\eta$ as
@@
$\Delta\phi(\Gamma)$ $=\int_{\gamma\circ\Gamma}d\varphi_{\mathbb{C}^{+}}$ (11)
The evaluation is illustrated in Fig. 2d) for the path $\Gamma$ from $\mathbf{w}_{1}$ to $\mathbf{w}^{\prime}$, shown in red. Mapping this path into $\mathbb{C}^{+}$, the path $\gamma\circ\Gamma$ is obtained. The unwrapped phase $\Delta\phi(\Gamma)$ is the angle spanned by $\gamma\circ\Gamma$ as seen from the origin [fulton97].
In the language of exterior calculus [weintraub14], the mapping $\gamma$ determines $d\varphi_{W^{+}}=\gamma^{\star}(d\varphi_{\mathbb{C}^{+}})$ as the pullback from $d\varphi_{\mathbb{C}^{+}}$. The pullback allows us to transfer evaluation of an infinitesimal phase change in the complex plane (through the one-form $d\varphi_{\mathbb{C}^{+}}$) to $W^{+}$ (through the one-form $d\varphi_{W^{+}}$).
2.2.3 Path dependence of phase unwrapping